OpenStudy (anonymous):

Hello, This is the latest that has stumped me... hang on I'm trying to see how to upload a little pic of the equation....

OpenStudy (anonymous):

Press attach file below

OpenStudy (jhannybean):

by "make \(L_1\) the subject" it is asking to solve for \(L_1\) correct?

OpenStudy (anonymous):

Cool thanks, I couldn't see how to post it straight up, but I see it now and have uploaded it :) thanks

OpenStudy (anonymous):

correct

OpenStudy (jhannybean):

\[\large \frac{1}{L}=\frac{1}{L_1}+\frac{1}{L_2+L_3}\]

jimthompson5910 (jim_thompson5910):

L1? It looks like L3

OpenStudy (anonymous):

yes thats right - what program do you use to be able to do that please??

OpenStudy (jhannybean):

Ahh, Sorry. \(L_3\)*

OpenStudy (jhannybean):

program??...We can simply solve for it.

OpenStudy (jhannybean):

First we need to find a common denominator between \(L~,~ L_1~,~ \text{and}~ L_2+L_3\).

OpenStudy (anonymous):

haha no i mean to be able to format the maths equation. not to solve the actual problem, jsut to be able to type in the equation like you have (rather than like what i had to do - go and scan my text book! ) There's gotta be a better way - looks like you know how to type it in using some program... you see what i was getting at?

OpenStudy (jhannybean):

Ah, this is just LaTex formatting. I simply typed in `\[\frac{1}{L}=\frac{1}{L_1}+\frac{1}{L_2+L_3}\]`

OpenStudy (jhannybean):

I've got to go right now, hopefully someone else can finish helping you with your question. Good luck!

OpenStudy (anonymous):

ok back to maths. So, to me the lowest common denominator would be L * L1 * L2 * L3 (im trying to say L times L1 times L2 times L3). Is that correct?

OpenStudy (anonymous):

ok thanks for the hint.

OpenStudy (mathstudent55):

Multiply each fraction by the least common denominator of all three fractions.

OpenStudy (mathstudent55):

\(LCD = LL_1(L_2 + L_3)\) Notice the plus sign between L2 and L3.

OpenStudy (anonymous):

Im trying it again with your advice now, thank you!

OpenStudy (mathstudent55):

\(\large LL_1(L_2 + L_3) \dfrac{1}{L}=LL_1(L_2 + L_3) \dfrac{1}{L_1}+LL_1(L_2 + L_3) \dfrac{1}{L_2+L_3}\)

OpenStudy (mathstudent55):

The purpose of multiplying each fraction by the LCD is to then eliminate all denominators.

OpenStudy (mathstudent55):

\( \cancel{L}L_1(L_2 + L_3) \dfrac{1}{\cancel{L}}=L\cancel{L_1}(L_2 + L_3) \dfrac{1}{\cancel{L_1}}+LL_1(\cancel{L_2 + L_3}) \dfrac{1}{\cancel{L_2+L_3}}\) \( L_1(L_2 + L_3) =L(L_2 + L_3)+LL_1\)

OpenStudy (anonymous):

thank you!! I'm following well so far. Im just expanding them out now to hopefully successfully isolate L3 .... :)

OpenStudy (mathstudent55):

\( L_1L_2 + L_1L_3 =LL_2 + LL_3+LL_1\) \(L_1L_3 - LL_3 = LL_1 + LL_2- L_1L_2\) \(L_3(L_1 - L) = LL_1 + LL_2- L_1L_2\) \(L_3 = \dfrac{LL_1 + LL_2- L_1L_2}{ L_1 - L } \)

OpenStudy (anonymous):

Argh you are amazing. I see :D THANK YOU! Far out thank so much !!!!!

OpenStudy (anonymous):

Hopefully this approach will help me with the others that Im getting wrong that are similar. It should :) Thanks so much !! If you have time and are up for it can I show you another one I am currently unable to get right? If not I will try and tackle ot with my new tools you have just given me :)

OpenStudy (anonymous):

Hi are you still there..? Im sorry but I double checked and the answer is different :( Its; LL1 L3 = ------ - L2 L1 - L To clarify, the answer in words is .... L3 = LL1 divided by L1 - L, minus L2

OpenStudy (mathstudent55):

\(L_3 = \dfrac{LL_1 + LL_2- L_1L_2}{ L_1 - L }\) \(L_3 = \dfrac{LL_1}{L_1 - L} + \dfrac{LL_2- L_1L_2}{ L_1 - L }\) \(L_3 = \dfrac{LL_1}{L_1 - L} - \dfrac{L_1L_2 - LL_2 }{ L_1 - L }\) \(L_3 = \dfrac{LL_1}{L_1 - L} - \dfrac{L_2(L_1- L) }{ L_1 - L }\) \(L_3 = \dfrac{LL_1}{L_1 - L} - \dfrac{L_2(\cancel{L_1- L}) }{\cancel{ L_1 - L} }\) \(L_3 = \dfrac{LL_1}{L_1 - L} - L_2 \)

OpenStudy (mathstudent55):

I started from our answer above and changed it into your answer. As you can see, both are the same just in different forms.

OpenStudy (anonymous):

oh boy - give me a sec haha :)

OpenStudy (mathstudent55):

gtg I'll read your comments tomorrow

OpenStudy (anonymous):

Yep, so they are (both the same just in different forms). Thanks very much for showing me all that, its helped me out a lot. THANKS :D

OpenStudy (mathstudent55):

You're very welcome.